A farm building costs $54,000 to build today and will earn after-tax net cash flows of $16,000 per year for five years. There is no salvage value on this building at the end of the five-year life, and the farmer’s cost of capital 9%.

a.   What is the net present value?

b.   What is the benefit-cost ratio of this building at a cost of capital of 9%?
Answer #1

fc =SUM(F1137:F1142) F1143 A B C 1135 Year 0 1136 1137 1138 1139 1140 1141 1142 1143 1144 1145 Cashflows PV@9% $ (54,000.00)- fx =SUM(F1137:F1142) F1143 А 1135 1136 Year Cashflows Pv@9% -54000 16000 16000 16000 16000 16000 1137 1138 1139 1140 1141 1